2x^2-4x+10=0 solve and complete the square
Answer by lwsshak3(6505)

2x^2-4x+10=0 solve and complete the square
2(x^2-2x+1)+10-2=0
2(x-1)^2+8=0
2(x-1)^2=-8
(x-1)^2=-4
take sqrt of both sides
x-1=√-4=2i
x=1±2i
no real solution